### v1

# Spin defects in hexagonal boron nitride as two-dimensional strain sensors

Abstract
Lattice deformation is a powerful way to engineer the properties of two-dimensional (2D) materials, making their precise measurement an important challenge for both fundamental science and technological applications. Here, we demonstrate that boron-vacancy (V) color centers in hexagonal boron nitride (hBN) enable quantitative strain sensing with sub-micrometer spatial resolution. Using this approach, we precisely quantify the strain-induced shift of the E2g Raman mode in a hBN flake under uniaxial stress, establishing V centers as a new tool for strain metrology in van der Waals heterostructures. Beyond strain sensing, our work also highlights the unique multimodal sensing functionalities offered by V centers, which will be valuable for future studies of strain-engineered 2D materials.
Strain engineering is a highly effective means of tuning the electronic, optical, and magnetic properties of two-dimensional (2D) materials and their associated van der Waals heterostructures Blundo et al. (2021); Qi et al. (2023); Bolotin et al. (2025). Over the past decade, a broad variety of methodologies have been developed to achieve precise control over lattice deformation in 2D materials, including deposition onto mechanically deformable substrates, nanoindentation, or bubble formation during crystal growth Blundo et al. (2021). However, fully exploiting strain engineering to tailor novel functionalities in 2D materials also requires parallel advances in techniques capable of providing independent, in situ measurements of lattice deformation with high spatial resolution. This need arises because the actual strain distribution in a 2D material often deviates from theoretical predictions owing to several effects that cannot be captured by conventional finite element simulations, such as layer slippage, substrate adhesion, edge reconstruction or interfacial and interlayer strain transmission Akinwande et al. (2017); Song et al. (2025).
In this Letter, we explore a novel approach to strain sensing in 2D materials that relies on a quantum sensing platform based on boron-vacancy (V) color centers in hexagonal boron nitride (hBN) Vaidya et al. (2023). This photoluminescent point defect features a spin-triplet ground state, whose electron spin resonance (ESR) frequencies can be interrogated via optically detected magnetic resonance (ODMR) spectroscopy under ambient conditions Gottscholl et al. (2020), even in the limit of ultra-thin hBN layers Durand et al. (2023). In addition, the ESR frequencies are highly sensitive to external perturbations such as magnetic and electric fields, pressure, temperature, or strain Liu et al. (2021); Gottscholl et al. (2021); Udvarhelyi et al. (2023); Yang et al. (2022); Lyu et al. (2022). These properties make V centers in hBN a promising candidate for the development of flexible, multifunctional sensors that can be readily integrated into van der Waals heterostructures for probing the physics of 2D materials with atomic-scale proximity Huang et al. (2022); Healey et al. (2023); Kumar et al. (2022); Mu et al. (2025); He et al. (2025). Here, we present the first experimental study of V centers under continuously tunable in-plane strain, exploiting the flexibility of thin hBN flakes as a host. Using an ensemble of V centers, we demonstrate quantitative strain sensing with sub-micrometer spatial resolution. We then use this capability to precisely calibrate the strain induced in a hBN flake under uniaxial stress, and we investigate the strain-induced shift of the E2g Raman mode. Despite the central role of hBN in 2D materials research, the impact of strain on its vibrational properties has received surprisingly little attention, with only a few studies published in the literature Cai et al. (2017); Androulidakis et al. (2018); Blundo et al. (2022). Our results provide a precise quantification of the strain dependence of the E2g Raman mode in hBN, revealing a shift of cm, which can serve as a reference for future studies of strain in van der Waals heterostructures. Beyond this result, our work highlights the potential of V centers as a robust tool for multimodal sensing in van der Waals heterostructures, with important implications for the study of strain-engineered 2D materials.
In our experiments, a flexible polyimide (PI) film was used as a strain-transmitting substrate. An elongated hBN flake with uniform thickness ( nm) was mechanically exfoliated from a commercial hBN crystal (HQ graphene) and transferred onto the PI substrate (see SI). A 100-nm thick Au coating was then deposited to clamp the hBN flake more firmly onto the strain-transmitting substrate, leaving only a m gap without top coating [see Fig. 1(b)]. The full stack was subsequently implanted with nitrogen at keV energy with a dose of ions/cm2 to create V centers in the hBN flake, and finally integrated into a custom-built strain cell composed of three piezoelectric actuators mounted in parallel (see SI) Hicks et al. (2014). Applying a positive voltage to the two outer piezoelectric actuators of the strain cell induces a stretching of the PI substrate, which effectively transfers tensile strain to the hBN flake, as schematically illustrated in Fig. 1(a).
The impact of strain on the optical and spin properties of V centers in hBN was investigated with a scanning confocal microscope operating under ambient conditions. A laser excitation at nm was focused onto the hBN flake integrated in the strain cell through a long-working-distance objective with a numerical aperture of . The PL signal was collected by the same objective, and guided either to a spectrometer or to a silicon avalanche photodiode operating in photon counting regime. A typical PL raster scan of a hBN flake is shown in Fig. 1(c). The PL signal features a broad spectral line in the near infrared that corresponds to the characteristic emission of V centers (see SI). ODMR spectroscopy was performed by applying a microwave excitation via an external loop antenna placed in close proximity to the sample. The same setup was also used to perform Raman spectroscopy of the hBN flakes.
We start by introducing the Hamiltonian describing the spin triplet ground state () of the V center in order to determine how its ESR frequencies evolve with strain. Neglecting the hyperfine interaction with neighboring nuclear spins, the spin Hamiltonian of the V center’s ground state reads
where are the dimensionless electron spin operators. The first term results from the spin-spin interaction between the two unpaired electrons of the defect, which leads to an axial zero-field splitting GHz between the and spin sublevels Gottscholl et al. (2020), where denotes the electron spin projection along the axis () of the hBN crystal. The second term arises from the interaction of the V center with a local electric field produced by surrounding charges in the hBN flake Gong et al. (2023); Durand et al. (2023); Udvarhelyi et al. (2023), which mixes the spin sublevels, yielding new eigenstates separated by an orthorhombic splitting [Fig. 2(a)]. This splitting increases with the density of charges in the hBN crystal and thus with the density of V centers. For the hBN flakes used in this work, MHz. The last term finally describes the spin-strain interaction, which is expressed as Udvarhelyi et al. (2023)
where are the strain coupling coefficients and the components of the strain tensor. Note that the out-of-plane strain components are not included because they have a negligible impact on the in-plane hBN structure that contains the electron spin density of the V center. The last two terms of the spin-strain Hamiltonian yield an additional orthorhombic splitting. However, ab initio calculations have shown that the corresponding strain coupling coefficients are very weak, such that the orthorhombic splitting is always dominated by the interaction with the local electric field in the strain range considered in this work () Udvarhelyi et al. (2023). The spin Hamiltonian can thus be simplified as
and the ESR frequencies of the V center are then given by , where .
First-principles calculations predict a strain coupling coefficient of MHz/% for a hBN monolayer, which increases to MHz/% for V centers embedded in thicker hBN flakes such as those used in this work Udvarhelyi et al. (2023). Tensile strain (taken positive by convention) thus leads to a reduction of the effective zero-field splitting parameter [Fig. 2(a)]. These theoretical predictions have been quantitatively confirmed by several recent experiments performed under hydrostatic stress conditions Mu et al. (2025); He et al. (2025); Yu et al. (2025). In the following, we employ the theoretically predicted coefficient to extract the local strain in the hBN flake from measurements of the ESR frequencies.
ODMR spectra recorded at the center of the hBN flake for different voltages applied to the strain cell are shown in Fig. 2(b). At zero voltage, we obtain GHz, corresponding to an initial tensile strain , which is likely transferred to the hBN flake when it is fixed on the strain cell. Raising the applied voltage results in an increased strain, which gradually shifts the magnetic resonances of the V center to lower frequencies. For voltages up to V, the strain transferred to the hBN flakes increases linearly [Fig. 2(c)] and the process is perfectly reversible, the initial strain being recovered when the voltage is swept back to V. In this voltage range, the hBN flake thus follows an elastic deformation behavior.
V spin sensors can also be used to infer the strain distribution within the hBN flake with a diffraction-limited spatial resolution (m). Figure 3 shows typical strain profiles recorded along the long axis () of the hBN flake for various voltages applied to the strain cell. These measurements reveal a slight strain gradient of about %/m, which emerges between the two sides of the strain cell. This gradient likely results from different adhesion of the hBN flake on the strain-transmitting substrate at the opposite edges of the gap
Several remarks have to be made at this stage. First, the orthorhombic splitting is not modified in the considered strain range [see inset in Fig. 2(c)], supporting our assumption that this parameter is dominated by the interaction with a local electric field. Second, the PL signal of V centers and the contrast of ODMR spectra are not significantly modified when tensile strain increases [see Fig. 2(b)]. These findings might seem in contradiction with recent studies on the optical and spin properties of V centers placed under hydrostatic pressure conditions in a diamond anvil cell (DAC), which reported a drastic reduction of both PL signal and ODMR contrast for pressures above a few GPa Mu et al. (2025); He et al. (2025); Yu et al. (2025). A key difference between experiments carried out within a DAC and with a strain cell is the presence or not of an out-of-plane () stress component. Although this component has a negligible impact on the ground state electron spin sublevels of the V centre Udvarhelyi et al. (2023), it likely modifies the structure of its excited states and the spin-dependent dynamics of optical cycles Ivády et al. (2020); Clua-Provost et al. (2024). Experiments performed under hydrostatic pressure in a DAC suggest that compressive out-of-plane stress results in a quenching of both PL signal and ODMR contrast owing to variations of non-radiative decay rates via a metastable singlet state during optical cycles Mu et al. (2025). These effects are likely not relevant to the present work because stress is applied in-plane. It would be interesting to study in future the regime of tensile, out-of-plane stress, which could lead to an enhancement of the PL signal, possibly providing new insight into the origin of the intrinsic dimness of the V center emission.
As a first application of V-based strain sensing, we now investigate the strain-induced variation of the E2g Raman mode in hBN, a topic scarcely documented in the literature despite the central role of hBN in 2D materials research. To this end, both Raman and ODMR spectra are sequentially measured at the center of the hBN flake while gradually increasing the voltage applied to the strain cell. The ODMR spectra provide a precise, in situ strain calibration at the exact location where Raman spectra are recorded. As shown in Fig. 4(a), the effective zero-field splitting parameter of the V center initially decreases linearly with the applied voltage, consistently with the previous experiment. It reaches a minimum value GHz at a voltage of V, corresponding to a maximum strain of . Interestingly, further elongation of the strain cell does not increase the strain in the hBN flake, but instead leads to its gradual release as the voltage rises from V to V, most likely due to layer slippage Androulidakis et al. (2018). When the voltage is swept back to V, the strain decreases linearly with a slope similar to that observed during the voltage ramp-up and reaches a compressive state at zero voltage [Fig. 4(a)]. This pronounced hysteretic behavior of strain in a hBN flake under uniaxial stress highlights the complexity of the mechanical response in layered 2D materials, which cannot be captured by standard numerical simulations.
With the strain precisely calibrated using V centers, we proceed with the analysis of the Raman spectra. At zero strain, the E2g Raman mode of hBN is observed at cm-1 and shifts linearly with strain, with a slope cm [Fig. 4(b)], corresponding to a Grneisen parameter . In addition to the overall shift of the Raman mode, the anisotropic strain distribution in the hBN flake is also expected to split the doubly-degenerate E2g mode owing to a lowered crystal symmetry Androulidakis et al. (2018). This splitting cannot be observed in our experiments, most likely because of the broad Raman linewidth ( cm-1) resulting from the implantation of the hBN flake with nitrogen atoms [see inset in Fig. 4(b)]. From similar measurements performed on other hBN flakes subjected to uniaxial stress, we infer an average shift of the E2g Raman mode cm and an average Grneisen parameter (see SI). These values, which are fully consistent with the few previous reports available in the literature Cai et al. (2017); Androulidakis et al. (2018); Blundo et al. (2022), will likely serve as a reference for future studies of strain in van der Waals heterostructures.
Beyond strain sensing, V centers in hBN enable simultaneous measurements of magnetic fields, a feature that could prove useful for exploring strain-controlled magnetic properties of 2D magnets Wang et al. (2020); Cenker et al. (2022). The impact of strain and magnetic field on the ESR frequencies can be easily disentangled by applying a bias magnetic field along the axis () of hBN. In the limit , the ESR frequencies are expressed as , where GHz/T is the electron spin gyromagnetic ratio. From an ODMR spectrum, strain can thus be inferred from the sum of the two ESR frequencies, while their difference directly yields the magnetic field variation. The shot-noise limited magnetic field sensitivity is given by , where is the contrast of the ODMR spectrum, its linewidth, and the rate of detected photons Dréau et al. (2011). Replacing by the strain coupling coefficient yields the sensitivity to strain. For the 20-nm-thick hBN flake studied in this work, we estimate a strain sensitivity together with a magnetic field sensitivity T/ (see SI). These performances could be further improved by at least one order of magnitude by optimizing the microwave excitation used for ODMR spectroscopy Wu et al. (2025) and by relying on isotopically purified hBN crystals Clua-Provost et al. (2023); Gong et al. (2024).
To conclude, we have demonstrated that V centers in hBN provide quantitative in situ strain measurements with sub-micrometer spatial resolution, and we used this capability to precisely quantify the strain-induced shift of the Raman mode. These results establish V centers as a new tool for strain metrology in van der Waals heterostructures. Compared to other strain sensing methods such as Raman spectroscopy, a key advantage of V centers in hBN lies in their ability to simultaneously probe several external perturbations with high sensitivity. This unique multimodal sensing functionality opens up exciting opportunities to investigate the impact of lattice deformations on the vibrational, magnetic, and electronic properties of 2D materials.
Acknowledgements - This work was supported by the French Agence Nationale de la Recherche through the project Qfoil (ANR-23-QUAC-0003), the program ESR/EquipEx+ 2DMAG (grant ANR-21-ESRE-0025), the project NanoX Q2D-SENS (grant ANR-17-EURE-0009) in the framework of the “Programme des Investissements d’Avenir”, and the Institute for Quantum Technologies in Occitanie. Z.M. acknowledges support from the National Science Foundation of China No. 52575671 and the Fundamental Research Funds for the Central Universities. Z.Z. acknowledges support from the National Natural Science Foundation of China (12404124), the Guangdong Basic and Applied Basic Research Foundation (2023A1515110419), the Pearl River Talent Recruitment Program (2023QN10X156), and the Open Project of State Key Laboratory of Silicon and Advanced Semiconductor Materials, Zhejiang University (SKL2025-12).